OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? (A)x = 6 (B)x = 4 (C)x = -3 (D)x = -2

OpenStudy (anonymous):

I think (D) is the answer

OpenStudy (aravindg):

think again ?

OpenStudy (anonymous):

wouldn't you use the midpoint formula

OpenStudy (aravindg):

\[midpoint =\dfrac{(x_1+x_2)}{2},\dfrac{(y_1+y_2)}{2}\]

OpenStudy (aravindg):

yes !

OpenStudy (anonymous):

okay

OpenStudy (aravindg):

substitute it here and solve : \[\large \dfrac{x-2}{2}=2 \]

OpenStudy (anonymous):

oh okay that is where I messed up

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

thank you

OpenStudy (aravindg):

yw :)